# Equivariant Hilbert and Ehrhart series under translative group actions

### v2

@@ 7. Alcoved polytopes
Let $W$ be an irreducible Weyl group of rank $d$ with associated root system $\Phi_{W}\subseteq\mathbb{R}^{d}$ and nondegenerate symmetric bilinear form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) on $\mathbb{R}^{d}$. We will follow [Bou02, Hum90] as standard references about Coxeter and Weyl groups.
@@
[…]
@@
This is justified, as $\mathscr{A}_{d}$ is the image of $\widetilde{\mathscr{A}}_{A_{d+1}}$ under the nonsingular linear transformation $f:V\to\mathbb{R}^{d}$ sending the basis $\alpha_{1},\ldots,\alpha_{d}$ of $V$ to the basis $\beta_{1},\ldots,\beta_{d}$ of $\mathbb{R}^{d}$ given by $\beta_{i}=({e_{i}-e_{i+1}})({{d+1}})$ for $11\leq i\leq d-$ 1 and $\beta_{d}={(e_{1}+e_{2}+\ldots+e_{d})}({{d+1}})$ (we use $e_{i}$ for the standard basis of $\mathbb{R}^{d}$, the image of $f$).
@@
(where the notation “$x\lessdot y$” means that $y$ covers $x$, i.e.: $x<y$ and there is no $z\in X$ such that $x<z<y$).
@@
As an example we compute the equivariant Ehrhart series of the Lipschitz polytope of [SS18, Example 2.6]. Consider, for a fixed $n>0$, the poset $T_{n}$ with $n$ elements of rank 00 (labeled $1$ through $n$) and one element of rank one (that we label by $n+1$) that is comparable with every other element. Note that this labeling is “natural” in the sense of [SS18].
[…]
@@
Let $G$ be the group of automorphisms of $T_{n}$, i.e., $G$ is the symmetric group $S_{n}$ acting by permutation of the elements $1,\ldots,n$ of rank 00. Sanyal and Stump describe the maximal simplices of the unimodular triangulation by alcoves as follows. Let $\tau\in S_{n+1}$ be any permutation of the set of elements of $T_{n}$. We call $\tau$ descent-compatible if, for every chain $a\lessdot b$ in $T_{n}$, the number of descents of the word $\tau(a)\tau(b)$ only depends on $b$: we call $d_{\tau}(b)$ this number. In our situation this means that either (1) $\tau(n+1)>\tau(i)$ for all $i\leq n$ or (2) $\tau(n+1)<\tau(i)$ for all $i\leq n$. We conclude that $\tau$ is descent-compatible if and only it is of one of two types.
@@
Now recall the group $G=S_{n}$ of permutations of $[n]\subseteq T_{n}$ and its action $\rho^{*}$ on $\mathbb{R}^{T_{n}}$ given by $\rho^{*}(g)(f)=f\circ g^{-1}$. Notice that $a$, $z$ as well as $e_{n+1}$ are fixed by every group element. Moreover, for every $I\subseteq[n]$ we have $\rho^{*}(g)(v_{I})=v_{g(I)}$ and so $\rho^{*}(g)\Delta_{I_{*},x}=\Delta_{g(I_{*}),x}$.

### v3

@@ 7. Alcoved polytopes
Let $W$ be an irreducible Weyl group of rank $d$ with associated root system $\Phi_{W}\subseteq\mathbb{R}^{d}$ and nondegenerate symmetric bilinear form $(\cdot,\cdot)$ on $\mathbb{R}^{d}$. We will follow [Bou02, Hum90] as standard references about Coxeter and Weyl groups.
@@
If the Weyl group $W$ is not specified, type $A$ is generally understood (e.g., as in [LP07]). In this case, say for type $A_{d+1}$, the construction of the arrangement $\widetilde{\mathscr{A}}_{A_{d+1}}=\{H_{k}^{\alpha}\}_{k\in\mathbb{Z},\alpha\in\Phi_{W}}$ described at the beginning of the section is an arrangement in the linear subspace $V\subseteq\mathbb{R}^{d+1}$ spanned by the vectors $\alpha_{i}:=\epsilon_{i}-\epsilon_{i+1}$, $1\leq i\leq d$ (we use $\epsilon_{i}$ for the standard basis vectors of $\mathbb{R}^{d+1}$ for better clarity), and the bilinear form $(\cdot\mid\cdot)$ is defined for every $v,w\in V$ as $(v\mid w)=\frac{v^{T}w}{d+1}$ [Bou02, Plate 1, p. 265].
@@
This is justified, as $\mathscr{A}_{d}$ is the image of $\widetilde{\mathscr{A}}_{A_{d+1}}$ under the nonsingular linear transformation $f:V\to\mathbb{R}^{d}$ sending the basis $\alpha_{1},\ldots,\alpha_{d}$ of $V$ to the basis $\beta_{1},\ldots,\beta_{d}$ of $\mathbb{R}^{d}$ given by $\beta_{i}=({e_{i}-e_{i+1}})({{d+1}})$ for $1\leq i\leq d-1$ and $\beta_{d}={(e_{1}+e_{2}+\ldots+e_{d})}({{d+1}})$ (we use $e_{i}$ for the standard basis of $\mathbb{R}^{d}$, the image of $f$).
@@
(where the notation $x\lessdot y$” means that $y$ covers $x$, i.e.: $x<y$ and there is no $z\in X$ such that $x<z<y$).
@@
As an example we compute the equivariant Ehrhart series of the Lipschitz polytope of [SS18, Example 2.6]. Consider, for a fixed $n>0$, the poset $T_{n}$ with $n$ elements of rank 0 (labeled $1$ through $n$) and one element of rank one (that we label by $n+1$) that is comparable with every other element. Note that this labeling is natural in the usual poset-theoretic sense.
n+1n+11122n1n-1nn
@@
Let $G$ be the group of automorphisms of $T_{n}$, i.e., $G$ is the symmetric group $S_{n}$ acting by permutation of the elements $1,\ldots,n$ of rank 0. Sanyal and Stump describe the maximal simplices of the unimodular triangulation by alcoves as follows. Let $\tau\in S_{n+1}$ be any permutation of the set of elements of $T_{n}$. We call $\tau$ descent-compatible if, for every chain $a\lessdot b$ in $T_{n}$, the number of descents of the word $\tau(a)\tau(b)$ only depends on $b$: we call $d_{\tau}(b)$ this number. In our situation this means that either (1) $\tau(n+1)>\tau(i)$ for all $i\leq n$ or (2) $\tau(n+1)<\tau(i)$ for all $i\leq n$. We conclude that $\tau$ is descent-compatible if and only it is of one of two types.
@@
Now recall the group $G=S_{n}$ of permutations of $n]\subseteq T_{n}$ and its action $\rho^{*}$ on $\mathbb{R}^{T_{n}}$ given by $\rho^{*}(g)(f)=f\circ g^{-1}$. Notice that $a$, $z$ as well as $e_{n+1}$ are fixed by every group element. Moreover, for every $I\subseteq[n]$ we have $\rho^{*}(g)(v_{I})=v_{g(I)}$ and so $\rho^{*}(g)\Delta_{I_{*},x}=\Delta_{g(I_{*}),x}$.